A set of 7 fair coins was tossed 259 times and the frequency of throws

observed were as follows:


Number of Heads : 0 1 2 3 4 5 6 7

Frequency of throws : 3 7 23 62 63 53 37 11


Fit a binomial distribution and find mean and standard deviation of fitted

distribution.


1
Expert's answer
2021-12-20T16:10:28-0500

mean:

"\mu=\sum x_if_i/N=4.07"


"p=\mu/n=4.07/7=0.58"

"q=1-p=0.42"


"N=259"

"n=7"

"f(x=k)=NC^k_np^kq^{n-k}"


"f(x=0)=259q^{7}=0.6\approx1"

"f(x=1)=259\cdot7pq^{6}=5.77\approx6"

"f(x=2)=259\cdot C^2_7p^2q^{5}=23.91\approx24"

"f(x=3)=259\cdot C^3_7p^3q^{4}=55.03\approx55"

"f(x=4)=259\cdot C^4_7p^4q^{3}=76"

"f(x=5)=259\cdot C^5_7p^5q^{2}=62.97\approx63"

"f(x=6)=259\cdot7p^6q=28.99\approx29"

"f(x=7)=259p^7=5.71\approx6"


for fitted data:

mean:

"\mu=\sum x_if_i/N=4.05"

standard deviation:

"\sigma=\sqrt{\frac{\sum f_ix_i-(\sum f_i x_i)^2/N}{N}}=1.32"